OpenStudy (anonymous):

m/4+25=34 6th grade XD

OpenStudy (anonymous):

subtract 25 from both sides then multiply by 4 to get rid of 1/4 in front of x

OpenStudy (anonymous):

1/4 m =9 m=36

OpenStudy (anonymous):

made a mistake sorry m/4+24=35 XD

OpenStudy (anonymous):

1/4 m = 11 m=44

OpenStudy (anonymous):

1/4 * 4 equals 1m=m whatever you do to one side you must do to the other

OpenStudy (anonymous):

to keep equation BALANCED

OpenStudy (anonymous):

thanks =D

OpenStudy (anonymous):

you're welcome!